OpenStudy (anonymous):

what is 15 divided by 2/5? please help!!!!

OpenStudy (anonymous):

\[15\div2/5\] \[15\times5/2=75/2=32.5\]

OpenStudy (anonymous):

@rrpking thank you but i think the answer is 49